# Beyond Complete Shapes: A Quantitative Evaluation of 3D Shape Matching Algorithms

### v1

@@ 11 Details about the BeCoS Benchmark and the Framework
11.1 Datasets used in BeCoS
As mentioned in the main paper we use seven different full datasets in our benchmark. We use each dataset entirely with two exceptions: For TOSCA we exclude the gorilla as it is excluded from most shape matching analysis due to its difficulty. The original dataset [36] of DT4D provides animated meshes for each category with a lot of categories. In the original dataset also exist shapes that are difficult to semantically match, e.g. a dragon. We use the same categories that were selected from DT4D-Matching [39]. This dataset provides their own cross-dataset correspondences but only for a version of the dataset where each mesh is around 8000 vertices. Since the shapes are only given re-meshed and not in the original resolution we use the original animated files but extract the same categories and frames from the animation as DT4D-Matching. This ensures that we have higher resolution input shapes and less error propagation along the correspondence paths as the shapes in the same category have the same triangulation. We end up with 39 categories for four-legged animals and eight categories of humanoid shapes.
12 Further Distrubtion Information of the Dataset
12.1 Hosting, Licences, Maintenance Plan & Long-Term Preservation
Our pipeline is based on existing 3D shape datasets. The original datasets have to be do downloaded and can only be used with the corresponding licences from the original datasets. We ensure this correct usage by letting the user download the data directly from the corresponding websites. Our code for generating the partial shapes will be available on GitHub with a CC BY-NC-SA 4.0 (open source for non-commercial use) licence.
Dataset License Website
FAUST [10] License link Project page
SCAPE [2] N.A. Project page
KIDS [45] N.A. Project page
SMAL [56] License link Project page
SHREC’20 [22] N.A. Project page
DT4D [36] License link Project page
TOSCA [13] N.A. We have permission
to distribute
Table 6: List of licenses of all datasets used in the BeCoS benchmark. SCAPE, KIDS, SHREC’20 and TOSCA do not have explicit licenses. We got permission from the original authors of TOSCA to distribute the assets on our GitHub repository.
Our main product is a publicly accessible framework, which we intend to permanently host on our GitHub repository. The biggest threat regarding accessibility is related to the BeCoS benchmark (which we consider as a specific instantiation of our universal framework) in case some of the involved 3D shape datasets become unavailable. However, given that the 3D shape datasets used in our framework are well-established, we anticipate their continued public availability. In the event that a dataset is taken down by the original authors, our framework’s functionality will remain intact for the remaining datasets, as we retain the generated shape correspondences.
@@ 12.1 Hosting, Licences, Maintenance Plan & Long-Term Preservation
DPFM [3] 8-20h 1.1s
SM-COMB [48] N/A 0.08h
GC-PPSM [24] N/A 2.5h
PFM [46] N/A 0.2h
ULRSSM [15] 16-86h 20s
Smooth Shells [25] N/A 0.2h
GeomFMaps [21] 3-7h 1.1s
Table 7: A break down of the compute time per method presented as baseline in our work. Axiomatic methods do not require any training time. We show training time for the entire training data and the evaluation time per shape pair.
13 Compute and Resources
We use an internal cluster to compute the given results with the different models. For Sm-Comb and Gc-Ppsm we use an Intel Xeon E5-2697 with 16 cores and up to 36GB of RAM per run. For the learning based methods we use an Intel Xeon E5-2697 with 16 cores and one Nvidia Titan GPU with 12G VRAM. As PFM’s precompiled code runs only on Windows, we use an Intel i7-14700KF with 20 cores and up to 32GB of RAM. We present the compute time per method used as baseline in our paper in Table 7.

### v2

@@ 10 Details about the BeCoS Benchmark and the Framework
As mentioned in the main paper, we use seven different full datasets in our benchmark. We use each dataset entirely with two exceptions: For TOSCA we exclude the gorilla as it is excluded from most shape matching analyses due to its difficulty. The original dataset [li20214dcomplete] of DT4D provides animated meshes for each category with a lot of categories. Some categories in the original dataset are difficult to semantically match to other categories, e.g. dragons. Therefore, we use the same categories that were selected from DT4D-Matching [magnet2022smooth]. This dataset provides its own cross-dataset correspondences but only for a version of the dataset where each mesh is around 8000 vertices. Since the shapes are only re-meshed and not in the original resolution, we use the original animated files but extract the same categories and frames from the animation as DT4D-Matching. This ensures that we have higher resolution input shapes and less error propagation along the correspondence paths, as the shapes in the same category have the same triangulation. We end up with 39 categories for four-legged animals and eight categories of humanoid shapes.
11 Hosting, Licences, Maintenance Plan & Long-Term Preservation of BeCoS
Our pipeline is based on existing 3D shape datasets. The original datasets have to be downloaded and can only be used with the corresponding licences from the original datasets. We ensure this correct usage by letting the user download the data directly from the corresponding websites. Our code for generating the (partial) shape pairs will be available on GitHub with a CC BY-NC-SA 4.0 (open source for non-commercial use) licence.
Our main product is a publicly accessible framework, which we intend to permanently host on our GitHub repository. Since the BeCoS benchmark is a specific instantiation of this universal framework, its accessibility depends on the availability of certain 3D shape datasets. The biggest risk regarding accessibility is therefore related to the BeCoS benchmark in case some of the involved 3D shape datasets become unavailable. However, given that the 3D shape datasets used in our framework are well-established, we anticipate their continued public availability. In the event that a dataset is taken down by the originators, our framework’s functionality will remain intact for the remaining datasets.
12 Compute and Resources
We use an internal cluster to compute the given results with the different models. For SM-COMB and GC-PPSM, we use an Intel Xeon E5-2697 with 16 cores and up to 36GB of RAM per run. For the learning based methods, we use an Intel Xeon E5-2697 with 16 cores and one Nvidia Titan GPU with 12G VRAM. As PFM’s precompiled code runs only on Windows, we use an Intel i7-14700KF with 20 cores and up to 32GB of RAM. We present the compute time per method used in our paper in Table. 7.
@@ 12 Compute and Resources
DPFM [attaiki2021dpfm] 8-20h 1.1s
SM-COMB [roetzer2022scalable] N/A 0.08h
GC-PPSM [ehm2024partial] N/A 2.5h
PFM [rodola2017partial] N/A 0.2h
ULRSSM [cao2023unsupervised] 16-86h 20s
Smooth Shells [eisenberger2020] N/A 0.2h
GeomFMaps [donati2020deepGeoMaps] 3-7h 1.1s
Table 7: A breakdown of the compute time per method presented as a baseline in our work. Axiomatic methods do not require any training time. We show the training time for the entire training data and the evaluation time per shape pair.
13 All Framework Options
In Table 8, we provide a full list of the general options available in our framework. We invite the reader to check our GitHub repository for further information.
Table 8: All options in the main configuration file. The default values of these options are stored on our framework code and allow the generation of the BeCoS benchmark.
Option Description
data_dir Path to data directory.
datasets True if the dataset is used.
combinations Type of shapes to generate, possible values:
human, four-legged, human_centaur,
four-legged_centaur, or all.
setting The setting to generate, possible values:
full_full/partial_full/partial_partial.
remesh True if the shapes should be re-meshed after
loading.
cam_pos_regime Specifies how far the pair of camera poses are
apart, possible values: high/medium/low.
store_vis If True, store visualization of each pair
of shapes.
show_output If True, an open3d visualization is shown during
generation.
original_settings If True, multiple assertions are activated to
ensure that BeCoS is generated.
use_precompute_remeshing If True, use the cached re-meshed shapes
if available.
update_precomputed_remeshed If True, update the caching directory for
re-meshed meshes.
use_precomputed_partial_raycasting If True, use the cached triangle IDs of each
partial shape if available.
update_precomputed_raycasting If True, update the caching directory for triangle
IDs.
one_axis_rotation If True, perform a random rotation around the
z-axis after generating the pair of shapes.
n_cam_pos The number of camera poses to sample to get
within the desired range of overlap.
min_overlap Minimum overalp between partial shapes desired.
max_overlap Maximum overalp between partial shapes desired.